Using Gauss’ Theorem calculate the flux of the vector field F ˆ
i ˆ
j kˆ = x + y + z r through
the surface of a cylinder of radius A and height H, which has its axis along the z-axis
and the base of the cylinder is on the xy-plane.
1
Expert's answer
2019-04-15T11:47:14-0400

Gauss' theorem states


"Flux=\int\overrightarrow{F}d\overrightarrow{A}=\int div\overrightarrow{F}dV"

"div\overrightarrow{F}={\partial F_x \over \partial x}+{\partial F_y \over \partial y}+{\partial F_z \over \partial z}"


"\overrightarrow{F}=x\overrightarrow{i}+y\overrightarrow{j}+z\overrightarrow{k}"

"div\overrightarrow{F}={\partial x \over \partial x}+{\partial y \over \partial y}+{\partial z \over \partial z}=1+1+1=3"

Hence


"Flux=\int\overrightarrow{F}d\overrightarrow{A}=\int div\overrightarrow{F}dV=\int 3 dV=3V=3\pi A^2H"